# Probabilistic Emulation of a Global Climate Model with Spherical DYffusion

### v1

@@ D.2.2 Climate variability
D.2.3 The impact of the ensemble size
Figure 10: We visualize the ensemble-mean RMSE, CRPS, and spread skill ratio as a function of the size, $m$, of the time-mean ensemble. For RMSE and CRPS we plot it relative to using a single time-mean prediction, i.e. $m=1$, of the same model. Error bars are computed by repeatedly subsampling different combinations of $m$-member ensembles given a superset of $M=25$ predictions.
In addition to the ensemble mean climate bias being generally significantly lower, in absolute terms than all the baselines, we find that our method’s ensemble performance also improves more dramatically than that of the probabilistic baselines as a function of the ensemble size relative to single-member performance. In  Figure 10, we visualize the relative improvement in terms of ensemble-mean RMSE and CRPS, as well as the spread skill ratio as a function of the ensemble size. The relative boost in performance in terms of ensemble-mean RMSE and CRPS is generally considerably larger for our method than for the probabilistic baselines. The effectiveness of our method’s ensemble follows closely the reference 10-member ensemble for surface pressure, but reveals a significant gap for the total water path field. Our method produces better dispersed time-mean ensembles compared to the baselines as evidenced by its spread skill ratio scores, which are closer to one, but still manifest a certain amount of underdispersiveness.
D.3 Weather forecasting
@@
Figure 11: Our method generates competitive probabilistic ensemble weather forecasts, a necessary but not sufficient prerequisite for achieving good climate simulations. We evaluate a 25-member ensemble as well as a single forecast for our method and DYffusion. We also report the deterministic forecasts of ACE.
@@
While we focus on climate time scales in this work, climate is formed by the statistics of weather, so it is important to verify that our method also generates reasonable forecasts of the weather simulated by the reference model. In Figure 11, we analyze the medium-range forecasting skill of our method and the baselines for lead times up to two weeks. Interestingly, ACE and DYffusion show persistent biases for the surface pressure field that are clearly visible from the first few days of forecasts already but do not seem to reflect on the RMSEs at weather time scales. Such persistent biases, however, may be magnified over longer simulations and could explain why the baselines have problems reproducing accurate long-term climate statistics. In terms of RMSE, the deterministic model ACE generally has a slight edge over our method and DYffusion, especially on lead times of less than a week. After that, the ensembles of our method and DYffusion perform best in terms of ensemble-mean RMSE. As expected, the ensemble mean significantly reduces the RMSE compared to using a single sample from our method or DYffusion, especially at longer lead times. The probabilistic metrics, CRPS, and spread / RMSE ratio show that our method’s and DYffusion’s ensemble perform quite similarly, even though they are based on completely different ML architectures. Both ensembles tend to be underdispersed (Spread / RMSE $<1$) on short time scales but quickly converge to a well-dispersed ensemble at longer lead times which persists for the whole 10-year climate simulations (not shown).
D.4 Weather vs. climate performance
@@
Figure 12: We visualize the performances of DYffusion and Spherical DYffusion (in different marker colors) at multiple checkpoint epochs and for multiple generated samples. We plot the 10-year time-mean RMSE of three example fields versus the time-step-wise near-surface temperature RMSE averaged out over the first 20 forecasts (5 days). The 5-day weather forecast performance shows no correlation with the long-term climate biases (indeed, there seems to exist an inverse correlation). This has important implications for practitioners, implying that optimizing for short-term forecasts alone – as is current practice for most ML-based weather forecasting models – may be suboptimal for attaining accurate climate simulations. We have verified that the behavior shown above holds for fields other than near-surface temperature too (not shown).

### v2

@@ E.1.2 Climate variability
E.2 Weather forecasting
@@
Figure 13: Comparison of medium-range weather forecasting skill between Spherical DYffusion (25-member ensemble and single forecast), DYffusion (25-member ensemble and single forecast), and ACE (single, deterministic forecast). Our method generates competitive probabilistic ensemble weather forecasts, a necessary but not sufficient prerequisite for achieving good climate simulations.
@@
While we focus on climate time scales in this work, climate is formed by the statistics of weather, so it is important to verify that our method also generates reasonable forecasts of the weather simulated by the reference model. In Figure 13, we analyze the medium-range forecasting skill of our method and the baselines for lead times up to two weeks. Interestingly, ACE and DYffusion show persistent biases for the surface pressure field that are clearly visible from the first few days of forecasts already but do not seem to reflect on the RMSEs at weather time scales. Such persistent biases, however, may be magnified over longer simulations and could explain why the baselines have problems reproducing accurate long-term climate statistics. In terms of RMSE, the deterministic model ACE generally has a slight edge over our method and DYffusion, especially on lead times of less than a week. After that, the ensembles of our method and DYffusion perform best in terms of ensemble-mean RMSE. As expected, the ensemble mean significantly reduces the RMSE compared to using a single sample from our method or DYffusion, especially at longer lead times. The ensemble metrics, CRPS, and spread / RMSE ratio show that our method’s and DYffusion’s ensemble perform quite similarly, even though they are based on completely different ML architectures. Both ensembles tend to be underdispersed (Spread / RMSE $<1$) on short time scales but quickly converge to a well-dispersed ensemble at longer lead times which persists for the whole 10-year climate simulations (not shown).
E.3 Weather vs. climate performance
@@
Figure 14: We visualize the performances of DYffusion and Spherical DYffusion (in different marker colors) at multiple checkpoint epochs and for multiple generated samples. We plot the 10-year time-mean RMSE (“climate skill”) of three example fields versus the time-step-wise near-surface temperature RMSE averaged out over the first 20 forecasts (5 days; “weather skill”). The 5-day weather forecast performance shows no correlation with the long-term climate biases (indeed, there seems to exist an inverse correlation). This has important implications for practitioners, implying that optimizing for short-term forecasts alone – as is current practice for most ML-based weather forecasting models – may be suboptimal for attaining accurate climate simulations. We have verified that the behavior shown above holds for fields other than near-surface temperature too (not shown).
@@
E.4 Qualitative samples
Figures 15, 17, and 17 compare near-surface air temperature, near-surface wind speed. and total water path between the FV3GFS validation simulation, two randomly selected 10-year trajectories generated by Spherical DYffusion, and the trajectory predicted by ACE. For both variables, we show the final ten snapshots of each simulation. The complete temporal evolutions of these simulations for near-surface wind speed and total water path can be viewed at https://youtu.be/7lHra7gBiBo and https://youtu.be/Hac_xGsJ1qY, respectively. The emulated fields demonstrate high realism, closely mimicking the patterns and variability observed in actual climate model outputs. This showcases Spherical DYffusion’s capability to generate plausible and physically consistent climate scenarios over decadal timescales.
Figure 15: We visualize the final 10 predictions from two random 10-year trajectory samples (i.e. the end of the ninth year) generated by Spherical DYffusion (middle rows) and ACE (bottom row). Here, we show the near-surface air temperature variable, $T_{k}$ for level $k=7$. It is important to note that at these extended time scales, simulated trajectories are expected to diverge significantly from one another for any given time step.
Figure 16: We visualize the final 10 predictions from two random 10-year trajectory samples generated by Spherical DYffusion (middle rows) and ACE (bottom row). Here, we show the derived near-surface wind speed variable, $\mathrm{WS}_{k}$ for level $k=7$. It is important to note that at these extended time scales, simulated trajectories are expected to diverge significantly from one another for any given time step. A video visualizing the full 10-year simulations is accessible at https://youtu.be/7lHra7gBiBo.
Figure 17: Same as Figure 17 but for the derived total water path variable, $\mathrm{TWP}$. A video visualizing the full 10-year simulations is accessible at https://youtu.be/Hac_xGsJ1qY.